# Post-detection inference for sequential changepoint localization

### v2

@@ B.2 Composite pre- and post-change
B.3 Experimental results
B.3.1 Known pre- and post-change (Setting I)
We consider Poisson the mean-change problem: $\text{Pois}(1)$ (pre-change) vs $\text{Pois}(2)$ (post-change) with the true (unknown) changepoint at $T=100,500$. We employ the CUSUM change detection method defined in (C.1) with thresholds at $A=1000,10000$. After detection, we employ Algorithm 2 to construct confidence sets for $T$. 5.1 and B.1 provide visualizations of the confidence sets and point estimates (2.7) of the changepoints across $5$ runs.
@@ B.3.1 Known pre- and post-change (Setting I)
Figure B.1: Setting I (known pre- and post-change Poisson distributions): The first $T-1$ observations are drawn from Pois$(1)$ and the rest from Pois$(2)$, with $T=100$ (top row) or $T=500$ (bottom row). Change is detected using CUSUM detector (C.1) with $A=10000$ (top row) or $A=1000$ (bottom row). The point estimate (2.7) is shown in vertical red dashed line and the confidence set (using Algorithm 2) is shown in red points. $B=N=100$, $\alpha=0.05,L=\infty$. Results of $5$ independent simulations are shown.
@@
We report the average size and the conditional and unconditional coverage rates across $100$ independent runs in Table B.2.
@@
Table B.2: Setting I: Average (of $100$ independent runs) size and coverage of the confidence sets using Algorithm 2 with $F_{0}=\text{Pois}(1),F_{1}=\text{Pois}(2),B=N=100$, $\alpha=0.05$, the average absolute deviation $\hat{T}$ (2.7) and average detection delay (given a true detection) of the detection algorithm used (C.1) with threshold $A$, while varying $T$ and $A$.
@@
For $A=10000$, conditional and unconditional coverages are equal, as the probability of false detection before the true change is negligible. However, for $T=500$ with $A=1000$, the presence of a significant number of false detections before the true change results in lower unconditional coverage. Nevertheless, the empirical conditional coverage rates meet the desired level, supporting our theory. We also report the average absolute deviation of $\hat{T}$ (defined in (2.7)) and delay (given no false detection) in the table.
B.3.2 Known pre-change and composite post-change (Setting II)
We perform experiments for Poisson rate-change scenarios having pre-change distribution as $\text{Pois}(1)$, and the true (unknown) post-change data distribution as $\text{Pois}(2)$, with the changepoint at $T=100,200,$ assuming the composite post-change model to be $\mathcal{P}_{1}=\{\text{Pois}(\theta):\theta>1.9\}$. For detection, we employ a weighted CUSUM detector (3.15) with thresholds at $A=1000,10000$. For weighted CUSUM, we employ a discrete weight distribution taking values $\theta_{1}=\theta,\theta_{2}=\theta+d,\theta_{3}=\theta+2d,\cdots,\theta_{10}=\theta+9d$, $d=0.2$, and $\theta=1.9$, with exponentially decaying weights $w_{i}=e^{-\frac{i-1}{2}}-e^{-\frac{i}{2}},\text{ for }i=1,\cdots,9,\text{ and }w_{10}=e^{-9/2}$, which appeared to us to be a sensible discretized exponential mixture over the alternative class. After detection, we construct confidence sets (as defined in (3.11)) for $T$.
We report the average size and the conditional and unconditional coverage across $100$ independent runs in (5.4). The empirical conditional coverage rates are always a bit higher than the theoretical bound $1-\alpha-\beta$ (which is 0.925 in our experiments). It is worth noting that both the lengths and the conditional coverage rates are higher compared to the known pre- and post-change settings (Table B.2). We also report the average absolute deviation of $\hat{T}$ (defined in (2.7)) and delay (given no false detection) in the table. We observe that the confidence sets in this setting tend to be a bit more conservative compared to those with known pre- and post-change distributions (Setting 1), exhibiting both higher coverage and larger size.
Table B.3: Setting II: Average (of $100$ independent runs) size and (conditional and unconditional) coverage rates of confidence sets (3.11) with $B=N=100$ and $\alpha=0.05,\beta=0.025,L=\infty$, the average absolute deviation $\hat{T}$ (3.5) and average detection delay (given no false detection) of the change detection method used (3.15) with threshold $A$, having pre- and post-change data-generating distributions as $\text{Pois}(1)$ and $\text{Pois}(2)$ respectively, while we vary the changepoint $T$ and $A$.

### v3

@@ C.2 Another concrete example: Poisson rate change problem
Appendix D Experiments for confidence intervals for pre/post change parameters
D.1 Setting II
Table D.1 contains conditional and marginal (empirical) coverages, (conditional) average of the sizes and of the confidence intervals (i.e., the upper and lower boundaries) (defined in (5.1)) for $\theta_{1}$ across 500500500 independent runs with $\alpha=\beta=\eta=0.05$. The data generation and change detection procedure remain the same as in the experiment described in Setting II of Section 6. Across all settings, the observed conditional coverage consistently exceeds the theoretical lower bound of 0.850.850.85, indicating that our method is reliable but conservative.
Table D.1: Confidence interval for $\theta_{1}$: known pre-change but unknown post-change. The pre-change parameter $\theta_{0}=0$ is known, the true (unknown) post-change parameter is $\theta_{1}=1$, while the algorithm only knows $\theta_{1}\in\Theta_{1}$, which we mention in the second column of the table. (Setting II)
T $\Theta_{1}$
Conditional coverage
Marginal coverage
Conditional Length Conditional CI
100 [0.75,)[0.75$\infty$)[ 0.75 , ∞ ) 0.98 0.98 1.37 [0.76, 2.13]
100 [0.9,)[0.9$\infty$)[ 0.9 , ∞ ) 0.98 0.96 1.23 [0.90, 2.16]
500 [0.75,)[0.75$\infty$)[ 0.75 , ∞ ) 0.99 0.93 1.48 [0.75, 2.23]
500 [0.9,)[0.9$\infty$)[ 0.9 , ∞ ) 0.98 0.91 1.31 [0.90, 2.21]
D.2 Setting III
Table D.2 contains conditional and marginal (empirical) coverages, (conditional) average of the sizes and of the confidence intervals (i.e., the upper and lower boundaries), which are defined in (5.2) and (5.1)), for $\theta_{0}$ and $\theta_{1}$ across 500500500 independent runs with $\alpha=\beta=\eta_{0}=\eta_{1}=0.05$. he data generation and change detection procedure remain the same as in the experiment described in Setting III of Section 6. Across all settings, the observed conditional coverage consistently exceeds the theoretical lower bound of 0.800.800.80.
Table D.2: Confidence intervals for $\theta_{0},\theta_{1}$ with target conditional coverages 0.850.850.85(Setting III): unknown pre- and post-change. The true (unknown) values of the parameters are $\theta_{0}=0$ and $\theta_{1}=1$, while the algorithm only knows $\theta_{i}\in\Theta_{i},i=0,1$, which are mentioned in the second and third columns of the table.
T
$\Theta_{0}$
$\Theta_{1}$
Cond. coverage
Marg. coverage
Cond. length Cond. CI
$\theta_{0}$ $\theta_{1}$ $\theta_{0}$ $\theta_{1}$ $\theta_{0}$ $\theta_{1}$ $\theta_{0}$ $\theta_{1}$
100 (,0.25]($\infty$,0.25]( ∞ , 0.25 ] [0.75,)[0.75$\infty$)[ 0.75 , ∞ ) 1.00 0.99 1.00 0.99 0.55 1.19 [-0.32, 0.23] [0.76, 1.95]
100 (,0.1]($\infty$,0.1]( ∞ , 0.1 ] [0.9,)[0.9$\infty$)[ 0.9 , ∞ ) 0.98 0.99 0.98 0.98 0.61 1.16 [-0.51, 0.10] [0.90, 2.06]
500 (,0.25]($\infty$,0.25]( ∞ , 0.25 ] [0.75,)[0.75$\infty$)[ 0.75 , ∞ ) 0.99 0.99 0.93 0.93 0.21 1.22 [-0.10, 0.11] [0.75, 1.97]
500 (,0.1]($\infty$,0.1]( ∞ , 0.1 ] [0.9,)[0.9$\infty$)[ 0.9 , ∞ ) 0.99 0.99 0.94 0.94 0.19 1.24 [-0.09, 0.10] [0.90, 2.14]
D.3 Comparison with [36]
Next, we compare our confidence interval for the post-change parameter with that proposed by [39] and discussed in [36]. We implement Wu’s interval as discussed in Chapter 4.3 of [36] (or Section of [39]), replacing $\theta$ by $\hat{\theta}$ in the bias and variance terms, which is given by $\hat{\theta}-\frac{4}{5\sqrt{d\hat{\theta}(\tau-\hat{\nu})}}\pm z_{\alpha/2}\frac{\sqrt{1-{13}/(16d\hat{\theta})}}{\sqrt{\tau-\hat{\nu}}}$, where $\hat{\theta}=\frac{T_{\tau}}{\tau-\hat{\nu}}$, $\tau=\tau^{\prime}_{\text{cusum}}$, and $z_{\alpha^{\prime}/2}$ is the $\alpha^{\prime}/2$ quantile of the standard normal distribution. We use CUSUM [18] (the same change detector as in the last experiment of Section 6), with $d=10$ for both methods. For our method, we assume the post-change space to be $\Theta_{1}=[0,\infty)$ with $\alpha=\beta=\eta=0.05$. We report the average size and the conditional and unconditional coverage rates of both methods across 500500500 independent runs in Table D.3, with a target conditional coverage of $\alpha^{\prime}=0.85$. The results show that the existing method [36], which is based on asymptotic approximations, undercovers $\theta_{1}$ (both conditionally and marginally). At the same time, our finite-sample confidence intervals are conservative — achieving higher-than-nominal (conditional) coverage at the expense of larger (conditional) length.
Table D.3: Comparison of the confidence sets for $\theta_{1}$ (target conditional coverage 0.850.850.85) with Wu [36] (Setting II)
$F_{0}$ $F_{1}$ $T$ Conditional Length Conditional Coverage Marginal Coverage
Ours Wu (2007) Ours Wu (2007) Ours Wu (2007)
${N}(-0.25,1)$ ${N}(0.25,1)$ 100 1.46 0.47 0.99 0.77 0.98 0.74
${N}(-0.3,1)$ ${N}(0.3,1)$ 100 1.43 0.51 0.98 0.76 0.98 0.75
${N}(-0.25,1)$ ${N}(0.25,1)$ 500 1.48 0.49 0.99 0.78 0.91 0.64
${N}(-0.3,1)$ ${N}(0.3,1)$ 500 1.49 0.51 0.99 0.79 0.91 0.71
Appendix E Additional experiments for composite pre/post change
In this section, we present experimental results for the alternative approach in which each set $\mathcal{S}_{t-1}$ and/or $\mathcal{S}_{\tau-t+1}^{\prime}$ is discretized using 50 equispaced grid points. Let, $\{\theta_{1}^{(0)},\cdots,\theta_{I}^{(0)}\}\subseteq\mathcal{S}_{t-1}$ and $\{\theta_{1}^{(1)},\cdots,\theta_{K}^{(1)}\}\subseteq\mathcal{S}_{\tau-t+1}^{\prime}$ be the grid points. Here, $I=K=50$.
Therefore, for $j\in\{1,\cdots,B\}$, we only need to generate $\{\epsilon_{n}^{j}\}_{n\in\mathbb{N}}\stackrel{{\scriptstyle iid}}{{\sim}}N(0,1)$ for $j=1,\cdots,B$. For $i\in[I],k\in[K]$, define
$Y_{n}^{j}(\theta_{i}^{(0)},\theta_{k}^{(1)})=\begin{cases}\theta_{i}^{(0)}+\epsilon_{n}^{j},\text{ if }n<t\\ \theta_{k}^{(1)}+\epsilon_{n}^{j},\text{ otherwise. }\end{cases}$ (E.1)
We run $\mathcal{A}$ on $\{Y_{n}^{j}(\theta_{i}^{(0)},\theta_{k}^{(1)})\}_{n}$ and denote $\tau_{i,k}^{j}$ as the detection time. We compute the test statistics based on $\tau_{i,k}^{j}$ and $Y_{1}^{j}(\theta_{i}^{(0)},\theta_{k}^{(1)}),\cdots,Y_{\tau_{i,k}^{j}}^{j}(\theta_{i}^{(0)},\theta_{k}^{(1)})$. Then, we consider the following confidence set:
$\mathcal{C}=\left\{t\in\{1,\cdots,\tau\}:M_{t}\leq\max_{i\in[I],k\in[K]}\operatorname{Quantile}(1-\alpha r_{t};M_{t},\{M_{t,L}^{j}(\theta_{i}^{(0)},\theta_{k}^{(1)})\}_{j=1}^{B})\right\}$ (E.2)
As a special case, when $\theta_{0}$ is known, then, $\theta_{0}=\theta_{1}^{(0)}=\cdots=\theta_{I}^{(0)}=\mathcal{S}_{t-1}.$
As shown in Table E.1, the confidence sets (E.2) have lengths and coverages that are comparable to those from the earlier method reported in Table 6.4 and Table 6.6. However, this grid-based approach incurs significantly higher computational cost. Therefore, for practical purposes, we recommend using the earlier, more computationally efficient method.
Table E.1: Composite pre/post change
T
Pre-change
Class
Post-change
Class
Conditional
coverage
Unconditional
coverage
Size Time (s)
100 $N(0,1)$$\{N(\mu,1):\mu>0.75\}$ 0.96 0.96 27.05 365.81
100 $N(0,1)$$\{N(\mu,1):\mu>0.9\}$ 0.95 0.95 24.62 360.48
100 $\{N(\mu,1):\mu\leq 0.25\}$$\{N(\mu,1):\mu>0.75\}$ 0.96 0.96 30.88 17572.26
100 $\{N(\mu,1):\mu\leq 0.1\}$$\{N(\mu,1):\mu>0.9\}$ 0.96 0.95 30.14 17092.11
Appendix F Simpler confidence sets for detection algorithms controlling PFA
In this section, we shall discuss how one can construct simpler confidence sets when the change detection algorithm $\mathcal{A}$ controls the probability of false alarm (PFA) at level $\delta\in(0,1)$. Importantly, this method allows us to avoid Assumption 3.3, and our results hold in the fully nonparametric case with arbitrary pre- and post-change classes, $\mathcal{P}_{0}$ and $\mathcal{P}_{1}$, respectively.
We consider the test statistic $M_{t}$ introduced in (3.4), and define the following confidence set for $T$:
$\mathcal{C}=\left\{t\in\mathbb{N}:t\leq\tau,M_{t}<\frac{2}{\alpha}\right\},$ (F.1)
which is a simplified version of Equation 3.8, obtained by directly replacing $\alpha r_{t}^{*}$ with $\alpha$. This avoids the need to compute $r_{t}^{*}$ (and hence $\tau_{1}^{*},\cdots,\tau_{N}^{*}$), making the method easier to implement and free of Assumption 3.3.
We now investigate the coverage properties of the above confidence set.
Theorem F.1.
Let $\mathcal{C}$ be the confidence set as defined in (F.1), for some $\alpha\in(0,1)$. If $\mathcal{A}$ controls PFA at level $\delta\in(0,1)$, then for any $F_{0}\in\mathcal{P}_{0}$ and $F_{1}\in\mathcal{P}_{1}$,
$\mathbb{P}_{F_{0},T,F_{1}}(T\in\mathcal{C})\geq 1-\alpha-\delta\leavevmode\nobreak\ \text{, and }\mathbb{P}_{F_{0},T,F_{1}}(T\in\mathcal{C}\mid\tau\geq T)\geq 1-\frac{\alpha}{1-\delta},$
where the latter guarantee holds if $\mathbb{\mathbb{P}}_{F_{0},\infty}(\tau\geq T)\neq 0,$ for all $F_{0}\in\mathcal{P}_{0}.$
Proof.
From Theorem 3.1, we have $\mathbb{P}_{F_{0},t,F_{1}}[M_{t}\geq\frac{2}{\alpha}]\leq\alpha$. Hence,
$\mathbb{P}_{F_{0},T,F_{1}}(M_{T}<2/\alpha)$
$=\mathbb{P}_{F_{0},T,F_{1}}(T\in\{t\in\mathbb{N}:M_{t}<2/\alpha\})$
$\leq\mathbb{P}_{F_{0},T,F_{1}}(T\in\{t\in\mathbb{N}:M_{t}<2/\alpha\},\tau\geq T)+\mathbb{\mathbb{P}}_{F_{0},T,F_{1}}(\tau<T)$
$=\mathbb{P}_{F_{0},T,F_{1}}(T\in\{t\in\mathbb{N}:t\leq\tau,M_{t}<2/\alpha\})+\mathbb{\mathbb{P}}_{F_{0},\infty}(\tau<T)$
$\leq\mathbb{P}_{F_{0},T,F_{1}}(T\in\mathcal{C})+\delta,$
where the last inequality follows from the fact that the PFA is less than or equal to $\delta$. Thus, we obtain
$\mathbb{P}_{F_{0},T,F_{1}}(T\in\mathcal{C})\geq 1-\alpha-\delta.$
Now,
$\mathbb{P}_{F_{0},T,F_{1}}(T\in\mathcal{C}\mid\tau\geq T)=\frac{\mathbb{P}_{F_{0},T,F_{1}}(T\in\mathcal{C})}{\mathbb{P}_{F_{0},T,F_{1}}(\tau\geq T)}=\frac{\mathbb{P}_{F_{0},T,F_{1}}(T\in\mathcal{C})}{\mathbb{P}_{F_{0},\infty}(\tau\geq T)}\geq 1-\frac{\alpha}{1-\delta},$
A similar simplification can be applied to the adaptive confidence sets introduced in Section 4 for the parametric setting. By replacing the $\alpha r_{t}^{*}$ term with $\alpha$ in Equations 4.1, 4.4 and 4.5, we obtain
$\mathcal{C}_{1}=\left\{t\in\mathbb{N}:t\leq\tau,M_{t}\leq\operatorname{Quantile}(1-{\alpha};M_{t},\{M_{t,L}^{j}\}_{j=1}^{B})\right\},$ (F.2)
for known pre- and post-change setting. Similarly, for the known pre-change and composite post-change setting,
$\mathcal{C}_{1}=\left\{t\in\mathbb{N}:t\leq\tau,M_{t}\leq\sup_{\theta\in\mathcal{S}_{\tau-t+1}}\operatorname{Quantile}(1-{\alpha};M_{t},\{M_{t,L}^{j}(\theta)\}_{j=1}^{B})\right\},$ (F.3)
And finally,
$\mathcal{C}_{1}=\left\{t\in\mathbb{N}:t\leq\tau,M_{t}\leq\sup_{(\theta,\theta^{\prime})\in\mathcal{\mathbf{S}}^{(t)}}\operatorname{Quantile}(1-{\alpha};M_{t},\{M_{t,L}^{j}(\theta,\theta^{\prime})\}_{j=1}^{B})\right\}$ (F.4)
for composite pre- and post-change setting.
For detection algorithms with PFA $\leq\delta,\leavevmode\nobreak\ \delta\in(0,1)$, these simpler forms retain the coverage guarantees — both conditional and unconditional — similar to what we have shown in the theorem above.
In the next section, we present the experimental results with the bounded PFA algorithms, along with some other experiments.
Appendix G Some additional experimental results
G.1 Experiments with dependent pre- and post-change data
[…]
We use the likelihood ratio-based cusum detector:
$\tau=\inf\left\{n\in\mathbb{N}:\max_{1\leq j\leq n}\left(\frac{p_{1}}{p_{0}}\right)^{n_{10}(j:n)}\left(\frac{1-p_{1}}{1-p_{0}}\right)^{n_{11}(j:n)}\geq A\right\},$ (G.1)
where $n_{11}(j:n)$ denotes the number of transitions from 1 to 1 and $n_{10}(j:n)$ denotes the number of transitions from 0 to 1 within $X_{j},\cdots,X_{n}$, and $A=1000$.
After detection, we compute the adaptive confidence sets (4.1) with the likelihood ratio-based test statistic (3.2) and point estimate as in (3.1). Note that our adaptive method (4.1) is the only one we know that is valid for dependent data. Table G.1 contains the results averaged across 500500500 runs.
Table G.1: Known pre- and post-change with dependency (Setting I)
T
Conditional
coverage
Marginal
coverage
Size
$|\hat{T}-T|$
Delay
100 0.99 0.98 36.12 4.35 28.99
500 0.98 0.98 35.83 4.42 27.36
G.2 Experiments with detection algorithms that control PFA for Gaussian mean change scenario
G.2.1 Known pre- and post-change (Setting A, PFA 0.001$\leq$ 0.001≤ 0.001)
We perform experiments for both Gaussian and Poisson mean-change scenarios having the true changepoint at $T=100,500$, using a likelihood ratio detector defined as
$\tau=\inf\left\{t\in\mathbb{N}:\prod_{i=1}^{t}\frac{f_{1}(X_{i})}{f_{0}(X_{i})}\geq A\right\},$ (G.2)
[…]
Table G.2: Known pre- and post-change distributions (Setting A, PFA 0.001$\leq$ 0.001≤ 0.001)
$T$ $F_{0}$ $F_{1}$ Conditional Coverage Conditional Size
$|\hat{T}-T|$
Delay
Universal Adaptive Universal Adaptive (Conditional) (Conditional)
100 $N(0,1)$ $N(1,1)$ 0.95 0.94 15.44 12.60 2.74 110.91
[…]
500 $N(0,1)$ $N(1,1)$ 0.97 0.97 15.92 12.65 2.24 514.12
[…]
G.2.2 Known pre- and unknown post-change (Setting B, PFA 0.001$\leq$ 0.001≤ 0.001)
$\tau=\inf\left\{t\in\mathbb{N}:\int_{\theta\in\Theta_{1}}\prod_{i=1}^{t}\frac{f_{\theta}(X_{i})}{f_{\theta_{0}}(X_{i})}dW(\theta)\geq A\right\}$ (G.3)
with thresholds at $A=1000$. The weights are chosen just as described in Table 6.2. Similarly, the PFA of this detector is at most $1/A$, i.e., 0.0010.0010.001 here. After detection, we construct our universal confidence sets (F.1) for $T$, with $\alpha=0.075$. The test statistic and the point estimate are the same as mentioned in Setting II of Section 6. We report the average size and the conditional and unconditional coverage rates across 500500500 independent runs in Table G.3. We also report the average absolute deviation of $\hat{T}$ in (B.2) and delay (given no false detection) in the table.
Table G.3: Known pre-change but unknown post-change (Setting B, PFA 0.001$\leq$ 0.001≤ 0.001)
T
Composite
post-change
Conditional
coverage
Unconditional
coverage
100 $\{N(\mu,1):\mu>0.75\}$ 0.96 0.95 23.82 3.68 118.76
100 $\{N(\mu,1):\mu>0.9\}$ 0.96 0.96 20.15 3.36 115.64
500 $\{N(\mu,1):\mu>0.75\}$ 0.97 0.97 23.15 3.18 530.83
500 $\{N(\mu,1):\mu>0.9\}$ 0.96 0.96 19.27 3.23 528.92
G.2.3 Composite pre- and post-change (Setting C, PFA 0.001$\leq$ 0.001≤ 0.001)
$\tau=\inf\left\{t\in\mathbb{N}:\int_{\theta\in\Theta_{1}}\prod_{i=1}^{t}\frac{f_{\theta}(X_{i})}{f^{*}_{\theta}(X_{i})}dW(\theta)\geq A\right\},$ (G.4)
[…]
Table G.4: Composite pre- and post-change (Setting C, PFA 0.001$\leq$ 0.001≤ 0.001)
T
Composite
pre-change
Composite
post-change
Conditional
coverage
Marginal
coverage
Size
$|\hat{T}-T|$
Delay
100 $\{N(\mu,1):\mu\leq 0.25\}$$\{N(\mu,1):\mu>0.75\}$ 0.97 0.97 26.02 3.97 125.64
100 $\{N(\mu,1):\mu\leq 0.1\}$$\{N(\mu,1):\mu>0.9\}$ 0.97 0.97 21.96 3.88 121.28
500 $\{N(\mu,1):\mu\leq 0.25\}$$\{N(\mu,1):\mu>0.75\}$ 0.97 0.97 25.18 3.82 544.91
500 $\{N(\mu,1):\mu\leq 0.1\}$$\{N(\mu,1):\mu>0.9\}$ 0.98 0.98 21.45 3.75 536.76
G.2.4 Huber-robust (nonparametric) pre- and post-change (Setting D, PFA 0.001$\leq$ 0.001≤ 0.001)
We perform experiments for the robust Gaussian mean-change problem under the total variation contamination model, with the true changepoint at $T=100$ and $T=500$. The pre- and post-change classes are total variation balls of radius $\epsilon$ around $N(\mu_{0},1)$ and $N(\mu_{1},1)$, respectively.
$\mathcal{P}_{i}=\{F\in\mathcal{M}:d_{TV}(F,N(\mu_{i},1))\leq\epsilon\};\leavevmode\nobreak\ \leavevmode\nobreak\ i=0,1.$
In our experiment, $\mu_{0}=0,\mu_{1}=1$ and $\epsilon=0.01,0.001$. The true (unknown) pre- and post-change distributions are $(1-\epsilon)\times N(0,1)+\epsilon\times\text{Cauchy}(-1,10)$ and $(1-\epsilon)\times N(1,1)+\epsilon\times\text{Cauchy}(-1,10)$ respectively. We use the Huber-robust tests by [23] as change-detectors:
$\tau=\inf\left\{t\in\mathbb{N}:\prod_{i=1}^{t}E_{i}\geq A\right\},$ (G.5)
[…]
After detection, we construct our universal confidence sets (F.1) for $T$, with $R_{n}^{(t)}=\prod_{i=t}^{n}E^{\prime}_{i}$, for $n\geq t$ $S_{n}^{(t)}=\prod_{i=n}^{t-1}E_{i}$, for $n<t$ and $\alpha=0.05$, where $E^{\prime}_{i}=\frac{\pi^{\prime}(X_{i})}{\mathbb{E}_{F_{1}}\pi^{\prime}(X_{i})+(c^{\prime\prime}-c^{\prime})\epsilon},$ $\pi^{\prime}(x)=\min\{\max\{\frac{f_{0}(x)}{f_{1}(x)},c^{\prime}\},c^{\prime\prime}\}$ and $E_{i}$ is as defined before. We report the average size and coverage rates across 200200200 independent runs in Table G.5. We also report the average absolute deviation of $\hat{T}$ and delay (given no false detection). The sizes and coverages are reasonable, despite the methods not knowing the data-generating distribution exactly.
Table G.5: Huber-robust (nonparametric) pre- and post-change (Setting D, PFA 0.001$\leq$ 0.001≤ 0.001)
T $\epsilon$
Conditional
coverage
Marginal
coverage
Size
$|\hat{T}-T|$
Delay
100 0.01 0.97 0.97 16.04 3.23 119.38
100 0.001 0.97 0.97 14.57 3.18 113.06
500 0.01 0.96 0.96 16.18 3.87 536.58
500 0.001 0.97 0.97 15.13 3.25 531.44
G.2.5 Comparison of experimental results with ARL and PFA controlling detection algorithms
G.3 Experimental results with Poisson distribution with detection algorithms that control ARL
G.3.1 Known pre- and post-change (Setting I)
[…]
@@ G.3.1 Known pre- and post-change (Setting I)
Figure G.6: Setting I (known pre- and post-change Poisson distributions): The first $T-1$ observations are drawn from Pois(1)(1)( 1 ) and the rest from Pois(2)(2)( 2 ), with $T=100,500$. The point estimate is shown in a vertical red dashed line, and the confidence set (using Algorithm 2) is shown in red points. $B=N=100$, $\alpha=0.05,L=\infty$. Results of 555 independent simulations are shown.
@@
We report the average size and the conditional and unconditional coverage rates across 500500500 independent runs in Table G.7.
@@
Table G.7: Setting I: Average (of 500500500 independent runs) size and coverage of the confidence sets using Algorithm 2 with $F_{0}=\text{Pois}(1),F_{1}=\text{Pois}(2),B=N=100$, $\alpha=0.05$, the average absolute deviation $\hat{T}$ and average detection delay (given a true detection) of the detection algorithm used (LABEL:eq:cusum) with threshold $A$, while varying $T$ and $A$.
@@
For $A=10000$, conditional and unconditional coverages are equal, as the probability of false detection before the true change is negligible. However, for $T=500$ with $A=1000$, the presence of a significant number of false detections before the true change results in lower unconditional coverage. Nevertheless, the empirical conditional coverage rates meet the desired level, supporting our theory. We also report the average absolute deviation of $\hat{T}$ and delay (given no false detection) in the table.
G.3.2 Known pre-change and composite post-change (Setting II)
[…]
$\tau_{\text{wcusum}}=\inf\left\{\tau\in\mathbb{N}:\max_{1\leq j\leq\tau}\int_{\theta\in\Theta_{1}}\prod_{i=j}^{\tau}\frac{f_{\theta}(X_{i})}{f_{\theta_{0}}(X_{i})}dW(\theta)\geq A\right\},$ (G.6)
[…]
We report the average size and the conditional and unconditional coverage across 500500500 independent runs in (G.8). The empirical conditional coverage rates are always a bit higher than the theoretical bound $1-\alpha-\beta$ (which is 0.925 in our experiments). It is worth noting that both the lengths and the conditional coverage rates are higher compared to the known pre- and post-change settings (Table G.7). We also report the average absolute deviation of $\hat{T}$ and delay (given no false detection) in the table. We observe that the confidence sets in this setting tend to be a bit more conservative compared to those with known pre- and post-change distributions (Setting 1), exhibiting both higher coverage and larger size.
[…]